# Optimal and Fair Encouragement Policy Evaluation and Learning

### v3

@@ C.2.3 Responder-dependent fairness measures
Importantly, the conditional exclusion restriction of Assumption 2 implies that responder status is independent of recommendation. Conditional on observables, whether a particular individual is a responder is independent of whether someone decides to treat them when recommended. In this way, we study responder status analogous to its use elsewhere in disparity assessment in algorithmic fairness [Imai et al., 2020, Kallus et al., 2019a]. Importantly, this assumption implies that the conditioning event (of being a responder) is therefore independent of the policy $\pi$, so it can be handled in the same framework. s
@@
Let $\mathcal{E}_{a}^{j}=\{A=a_{j}\},\mathcal{E}_{\circ}=\{\text{True}\},$ and we express Equation 4 as a set of equality constraints of the above moment $h_{a}(\pi)=h_{\circ}(\pi),$ leading to pairs of inequality constraints,
@@ C.2.4 Best-response oracles
[…]
@@
$\operatorname{BEST}_{\boldsymbol{\lambda}}(Q)$ is the best response of the $\Lambda$ player. It can be chosen to be either 00 or put all the mass on the most violated constraint. Let ${{\gamma}}(Q):={M}{{h}}(Q)$ denote the constraint values, then $\operatorname{BEST}_{\boldsymbol{\lambda}}(Q)$ returns
@@ C.3 Proofs
where the last line follows by the conditional exclusion restriction (Assumption 2) and consistency (Assumption 1).
Proof of Proposition 7.
@@
$\frac{\sum_{r}\mathbb{E}[\pi_{r}(X)p_{1\mid r}(X,A)(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a])}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a]}$ by Proposition 1
@@ D.1 Proofs for generalization under unconstrained policies
Assume the nuisance models $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0},e_{r}(X)]^{\top},\eta\in\mathcal{F}_{\eta}$ are consistent and well-specified with finite VC-dimension $v_{\eta}$ over the product function class $H$. We provide a proof for the general case, including doubly-robust estimators, which applies to the statement of Proposition 8 by taking $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0}].$
@@
Applying Equation 5 to each of the component classes $\mathcal{F}_{L_{1}}(O_{1:n}),\mathcal{F}_{L_{2}}(O_{1:n}),\mathcal{F}_{L_{3}}(O_{1:n}),$ and Lipschitz contraction [Bartlett and Mendelson, 2002, Thm. 12.4] of the positive part function $\mathcal{F}_{L_{+}}$, we obtain the bound
@@
Proof of Proposition 9.
The characterization follows by strong duality in infinite-dimensional linear programming [Shapiro, 2001]. Strict feasibility can be satisfied by, e.g. solving LABEL:eqn-feas-constr to set ranges for $\epsilon$. ∎
@@ D.2 Proofs for robust characterization
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[c_{rt}(Y(t))\mid X]P(T(r)=t\mid R=r,X)]$ Assumption 2 (ER) (6)
@@
[…]
Proof of proposition 5.
@@ D.3.1 Preliminaries: results from other works used without proof
Let $\xi:=\max_{h}\|{M}\hat{{\mu}}(h)-\hat{{c}}\|_{\infty}$. Suppose Assumption 7 holds for $C^{\prime}\geq 2C+2+\sqrt{\ln(4/\delta)/2}$, where $\delta>0$. Let $Q^{\star}$ minimize $V(Q)$ subject to ${M}{h}(Q)\leq{c}$. Then Algorithm 1 with $\nu\propto n^{-\alpha},B\propto n^{\alpha}$ and $\omega\propto\xi^{-2}n^{-2\alpha}$ terminates in $O\left(\xi^{2}n^{4\alpha}\ln|\mathcal{K}|\right)$ iterations and returns $\hat{Q}$. If $np_{j}^{\star}\geq 8\log(2/\delta)$ for all $j$, then with probability at least $1-(|\mathcal{J}|+1)\delta$ then for all $k$, $\hat{Q}$ satisfies:
@@ D.4 Adapted lemmas
D.5 Proof of Theorem 3
Proof of Theorem 3.
@@ D.5 Proof of Theorem 3
$\Pi_{2}\left(\epsilon_{n}\right)$ =$=\left\{\pi\in{\Pi}:\mathbb{E}_{n_{1}}[v(\pi;\eta_{0})-v(\hat{Q}_{1};\eta_{0})]\leq\epsilon_{n},\mathbb{E}_{n_{1}}\left[g_{j}(O;\pi,\eta_{0})-g_{j}\left(O;\hat{Q}_{1},\eta_{0}\right)\mid\mathcal{E}_{j}\right]\leq\epsilon_{n},j\in\hat{\mathcal{I}}_{1}\right\}$
$\mathcal{Q}_{2}\left(\epsilon_{n}\right)$ =$=\left\{Q\in\Delta\left({\Pi_{2}}(\epsilon_{n})\right)\right\}$
$\mathcal{Q}^{*}\left(\epsilon_{n}\right)$ =$=\left\{Q\in\Delta({\Pi}):\mathbb{E}[(v(Q;\eta_{0})-v(Q^{*};\eta_{0})]\leq\epsilon_{n},\mathbb{E}[g_{j}(O;Q,\eta_{0})\mid\mathcal{E}_{j}]-\mathbb{E}[g_{j}(O;Q^{*},\eta_{0})\mid\mathcal{E}_{j}]\leq\epsilon_{n}\right\}$
@@
Invoking Theorem 5 on the output of the first stage of the algorithm, yields that with probability $1-\frac{\delta}{6}$ over the randomness in $\mathcal{D}_{1},$ by choice of $\epsilon_{n}$ = $\bar{O}(n^{-\alpha}))$,
@@
Step 2: Again invoking Theorem 5, this time on the output of the second stage of the algorithm with function space $\Pi_{2}$ (hence implicitly $\mathcal{Q}_{2}$), and conditioning on the “good event" that $Q^{*}\in\mathcal{Q}_{2}$, we obtain the bound that with probability $\geq 1-\delta/3$ over the randomness of the second sample $\mathcal{D}_{2},$
@@
Note that for $Q\in\mathcal{Q}$, w.h.p. $61-\delta/$ 6 over the first sample, we have that
@@ D.6 Proofs of auxiliary/adapted lemmas
$\leq L(Q^{*},\lambda^{*};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$ by Lemma 4
@@
$\leq L(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\epsilon_{n}+\nu+\chi_{n,\delta}^{2}$ assuming $\text{assuming }\epsilon_{n}\geq\chi_{n,\delta}^{2}$
@@
Applying Lemma 4 gives;
@@
Following the rest of the argumentation in [Chernozhukov et al., 2019, Lemma 9] from here onwards gives the result, i.e. studying the case of estimated nuisances with our Lemma 5 and Lemma 4. ∎
@@ Appendix E Additional case study
E.1 Oregon Health Insurance Study
@@
  • $A$: race (non-white/white), or gender (female/male)
@@
Figure 8: Policy value $V(\pi^{\lambda})$, treatment value $\mathbb{E}[T(\pi^{\lambda})\mid A=a],$ for $A$ = race, gender.
In Figure 8 we plot descriptive statistics. We include histograms of the treatment responsivity lifts $p_{1\mid 1a}(x,a)-p_{1\mid 0a}(x,a).$ We see some differences in distributions of responsivity by gender and race. We then regress treatment responsivity on the outcome-model estimate of $\tau$. We find substantially more heterogeneity in treatment responsivity by race than by gender: whites are substantially more likely to take up insurance when made eligible, conditional on the same expected treatment effect heterogeneity in increase in healthcare utilization. (This is broadly consistent with health policy discussions regarding mistrust of the healthcare system).
Next we consider imposing treatment parity constraints on an unconstrained optimal policy (defined on these estimates). In Figure 8 we display the objective value, and $\mathbb{E}[T(\pi)\mid A=a]$, for gender and race, respectively, enumerated over values of the constraint. We use costs of $2$ for the number of doctors visits and $1$ for enrollment in Medicaid (so $\mathbb{E}[T(\pi)\mid A=a]$ is on the scale of probability of enrollment). These costs were chosen arbitrarily. Finding optimal policies that improve disparities in group-conditional access can be done with relatively little impact to the overall objective value. These group-conditional access disparities can be reduced from $4$ percentage points ($0.04$) for gender and about $6$ percentage points $(0.06)$ for race at a cost of $0.01$ or $0.02$ in objective value (twice the number of doctors’ visits). On the other hand, relative improvements/compromises in access value for the “advantaged group" show different tradeoffs. Plotting the tradeoff curve for race shows that, consistent with the large differences in treatment responsivity we see for whites, improving access for blacks. Looking at this disparity curve given $\lambda$ however, we can also see that small values of $\lambda$ can have relatively large improvements in access for blacks before these improvements saturate, and larger $\lambda$ values lead to smaller increases in access for blacks vs. larger decreases in access for whites.
E.2 Additional Discussion, PSA-DMF case study

### v4

@@ C.2.3 Responder-dependent fairness measures
Importantly, the conditional exclusion restriction of Assumption˜2 implies that responder status is independent of recommendation. Conditional on observables, whether a particular individual is a responder is independent of whether someone decides to treat them when recommended. In this way, we study responder status analogous to its use elsewhere in disparity assessment in algorithmic fairness [Imai et al., 2020, Kallus et al., 2019a]. Importantly, this assumption implies that the conditioning event (of being a responder) is therefore independent of the policy $\pi$, so it can be handled in the same framework. s
@@
Let $\mathcal{E}_{a}^{j}=\{A=a_{j}\},\mathcal{E}_{\circ}=\{\text{True}\},$ and we express Equation˜4 as a set of equality constraints of the above moment $h_{a}(\pi)=h_{\circ}(\pi),$ leading to pairs of inequality constraints,
@@ C.2.4 Best-response oracles
$\begin{aligned} L(\pi,{\lambda})&=\hat{V}(\pi)+{\lambda}^{\top}({M}\hat{{h}}(\pi)-\hat{{d}})\\ &=\hat{V}(\pi)-{\lambda}^{\top}\hat{{d}}+\sum_{k,j}\frac{M_{k,j}\lambda_{k}}{p_{j}}{\mathbb{E}}_{n}\left[g_{j}(O,\pi){1}\left\{O\in\mathcal{E}_{j}\right\}\right].\end{aligned}$
@@
$\operatorname{BEST}_{\boldsymbol{\lambda}}(Q)$ is the best response of the $\Lambda$ player. It can be chosen to be either 0 or put all the mass on the most violated constraint. Let ${{\gamma}}(Q):={M}{{h}}(Q)$ denote the constraint values, then $\operatorname{BEST}_{\boldsymbol{\lambda}}(Q)$ returns
@@ C.3 Proofs
where the last line follows by the conditional exclusion restriction (Assumption˜2) and consistency (Assumption˜1).
Proof of Proposition˜8.
@@
$\frac{\sum_{r}\mathbb{E}[\pi_{r}(X)p_{1\mid r}(X,A)(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a])}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a]}$ by Proposition˜2
@@ D.1 Proofs for generalization under unconstrained policies
Assume the nuisance models $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0},e_{r}(X)]^{\top},\eta\in\mathcal{F}_{\eta}$ are consistent and well-specified with finite VC-dimension $v_{\eta}$ over the product function class $H$. We provide a proof for the general case, including doubly-robust estimators, which applies to the statement of Proposition˜9 by taking $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0}].$
@@
Applying Equation˜5 to each of the component classes $\mathcal{F}_{L_{1}}(O_{1:n}),\mathcal{F}_{L_{2}}(O_{1:n}),\mathcal{F}_{L_{3}}(O_{1:n}),$ and Lipschitz contraction [Bartlett and Mendelson, 2002, Thm. 12.4] of the positive part function $\mathcal{F}_{L_{+}}$, we obtain the bound
@@
Proof of Proposition˜10.
The characterization follows by strong duality in infinite-dimensional linear programming [Shapiro, 2001]. Strict feasibility can be satisfied by, e.g. solving eq.˜2 to set ranges for $\epsilon$. ∎
@@ D.2 Proofs for robust characterization
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[c_{rt}(Y(t))\mid X]P(T(r)=t\mid R=r,X)]$ Assumption˜2 (ER) (6)
@@
$\text{ where }c^{*}_{rt}=\begin{cases}\overline{B}\mathbb{I}\left[t=1\right]+\underline{B}\mathbb{I}\left[t=0\right]&\text{ if }\mathbb{E}[Y\pi_{r}(X)\mid T=t]\geq 0\\ \overline{B}\mathbb{I}\left[t=0\right]+\underline{B}\mathbb{I}\left[t=1\right]&\text{ if }\mathbb{E}[Y\pi_{r}(X)\mid T=t]<0\end{cases}$
Proof of proposition˜6.
@@ D.3.1 Preliminaries: results from other works used without proof
Let $\xi:=\max_{h}\|{M}\hat{{\mu}}(h)-\hat{{c}}\|_{\infty}$. Suppose Assumption˜7 holds for $C^{\prime}\geq 2C+2+\sqrt{\ln(4/\delta)/2}$, where $\delta>0$. Let $Q^{\star}$ minimize $V(Q)$ subject to ${M}{h}(Q)\leq{c}$. Then Algorithm 1 with $\nu\propto n^{-\alpha},B\propto n^{\alpha}$ and $\omega\propto\xi^{-2}n^{-2\alpha}$ terminates in $O\left(\xi^{2}n^{4\alpha}\ln|\mathcal{K}|\right)$ iterations and returns $\hat{Q}$. If $np_{j}^{\star}\geq 8\log(2/\delta)$ for all $j$, then with probability at least $1-(|\mathcal{J}|+1)\delta$ then for all $k$, $\hat{Q}$ satisfies:
@@ D.4 Adapted lemmas
D.5 Proof of Theorem˜4
Proof of Theorem˜4.
@@ D.5 Proof of Theorem˜4
$\Pi_{2}\left(\epsilon_{n}\right)$ $=\left\{\pi\in{\Pi}:\mathbb{E}_{n_{1}}[v(\pi;\eta_{0})-v(\hat{Q}_{1};\eta_{0})]\leq\epsilon_{n},\mathbb{E}_{n_{1}}\left[g_{j}(O;\pi,\eta_{0})-g_{j}\left(O;\hat{Q}_{1},\eta_{0}\right)\mid\mathcal{E}_{j}\right]\leq\epsilon_{n},j\in\hat{\mathcal{I}}_{1}\right\}$
$\mathcal{Q}_{2}\left(\epsilon_{n}\right)$ $=\left\{Q\in\Delta\left({\Pi_{2}}(\epsilon_{n})\right)\right\}$
$\mathcal{Q}^{*}\left(\epsilon_{n}\right)$ $=\left\{Q\in\Delta({\Pi}):\mathbb{E}[(v(Q;\eta_{0})-v(Q^{*};\eta_{0})]\leq\epsilon_{n},\mathbb{E}[g_{j}(O;Q,\eta_{0})\mid\mathcal{E}_{j}]-\mathbb{E}[g_{j}(O;Q^{*},\eta_{0})\mid\mathcal{E}_{j}]\leq\epsilon_{n}\right\}$
@@
Invoking Theorem˜6 on the output of the first stage of the algorithm, yields that with probability $1-\frac{\delta}{6}$ over the randomness in $\mathcal{D}_{1},$ by choice of $\epsilon_{n}$ = $\bar{O}(n^{-\alpha}))$,
@@
Step 2: Again invoking Theorem˜6, this time on the output of the second stage of the algorithm with function space $\Pi_{2}$ (hence implicitly $\mathcal{Q}_{2}$), and conditioning on the “good event" that $Q^{*}\in\mathcal{Q}_{2}$, we obtain the bound that with probability $\geq 1-\delta/3$ over the randomness of the second sample $\mathcal{D}_{2},$
@@
Note that for $Q\in\mathcal{Q}$, w.h.p. $1-\delta/6$ over the first sample, we have that
@@ D.6 Proofs of auxiliary/adapted lemmas
$\leq L(Q^{*},\lambda^{*};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$ by Lemma˜4
@@
$\leq L(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\epsilon_{n}+\nu+\chi_{n,\delta}^{2}$ $\text{assuming }\epsilon_{n}\geq\chi_{n,\delta}^{2}$
@@
Applying Lemma˜4 gives;
@@
Following the rest of the argumentation in [Chernozhukov et al., 2019, Lemma 9] from here onwards gives the result, i.e. studying the case of estimated nuisances with our Lemma˜5 and Lemma˜4. ∎
@@ Appendix E Additional case study
E.1 Text message reminders
Table 3: Regression Results, $E[R\mid T_{0},X]$
Coefficient Std. Error t-statistic p-value
Intercept 0.000
HH Size 0.598*** (0.055) 10.826 0.000
Phone interview 0.730*** (0.049) 14.952 0.000
Female 0.252*** (0.044) 5.667 0.000
Nonwhite 0.236*** (0.047) 5.014 0.000
Age 0.032*** (0.002) 17.943 0.000
Citizen Status -0.290 (1.001) -0.290 0.772
First SNAP year -0.001 (0.024) -0.022 0.983
Any kids 0.429*** (0.096) 4.470 0.000
ESL 0.384 (1.001) 0.384 0.701
HH receives max amt -0.185*** (0.050) -3.718 0.000
No earnings prev quarter 0.001** (0.000) 2.118 0.034
Years since first SNAP 0.012 (0.022) 0.563 0.573
Interview week in month -0.290 (1.001) -0.290 0.772
Interview Day -0.020*** (0.003) -7.244 0.000
English Lang. Int. -0.384 (1.001) -0.383 0.701
Spanish Lang. Int. 0.151 (0.119) 1.262 0.207
reminder 0.034 (0.047) 0.720 0.472
Note: *** p<0.01, ** p<0.05, * p<0.1
Table 4: Regression Results, Interacted logistic regression $P(T\mid X,R,R\times X)$
Coefficient Std. Error t-statistic p-value
Intercept 0.000
HH Size 0.128*** (0.037) 3.516 0.000
Phone interview 0.141** (0.059) 2.380 0.017
Female 0.051 (0.048) 1.063 0.288
Nonwhite 0.048 (0.057) 0.845 0.398
Age 0.045*** (0.002) 22.244 0.000
Citizen Status -0.059 (1.001) -0.059 0.953
First SNAP year -0.000 (0.031) -0.016 0.987
Any kids 0.089 (0.082) 1.086 0.277
ESL 0.078 (1.001) 0.078 0.938
HH receives max amt -0.039 (0.052) -0.756 0.450
No earnings prev quarter -0.005*** (0.001) -8.802 0.000
Years since first SNAP 0.088*** (0.029) 3.014 0.003
Interview week in month -0.059 (1.001) -0.059 0.953
Interview Day -0.009*** (0.003) -3.137 0.002
English Lang. Int. -0.078 (1.001) -0.078 0.938
Spanish Lang. Int. 0.031 (0.089) 0.348 0.728
reminder 0.000 (1.414) 0.000 1.000
R $\times$ HH Size 0.033 (0.079) 0.411 0.681
R $\times$ Phone interview 0.031 (0.118) 0.261 0.794
R $\times$ Female 0.012 (0.093) 0.125 0.901
R $\times$ Nonwhite 0.010 (0.109) 0.090 0.928
R $\times$ Age -0.067*** (0.004) -16.384 0.000
R $\times$ Citizen Status -0.015 (1.003) -0.015 0.988
R $\times$ First SNAP year 0.001 (0.001) 1.058 0.290
R $\times$ Any kids 0.022 (0.186) 0.117 0.907
R $\times$ ESL 0.017 (1.006) 0.017 0.986
R $\times$ HH receives max amt -0.007 (0.106) -0.068 0.946
R $\times$ No earnings prev quarter 0.009*** (0.001) 8.733 0.000
R $\times$ Years since first SNAP 0.045*** (0.014) 3.237 0.001
R $\times$ Interview week in month -0.015 (1.003) -0.015 0.988
R $\times$ Interview Day 0.003 (0.006) 0.516 0.606
R $\times$ English Lang. Int. -0.017 (1.006) -0.017 0.986
R $\times$ Spanish Lang. Int. 0.014 (0.231) 0.060 0.952
Note: *** p<0.01, ** p<0.05, * p<0.1
E.2 Oregon Health Insurance Study
@@
  • AA: race (non-white/white), or gender (female/male)
@@
Figure 8: Policy value $V(\pi^{\lambda})$, treatment value $\mathbb{E}[T(\pi^{\lambda})\mid A=a],$ for AA = race, gender.
In Figure˜8 we plot descriptive statistics. We include histograms of the treatment responsivity lifts $p_{1\mid 1a}(x,a)-p_{1\mid 0a}(x,a).$ We see some differences in distributions of responsivity by gender and race. We then regress treatment responsivity on the outcome-model estimate of $\tau$. We find substantially more heterogeneity in treatment responsivity by race than by gender: whites are substantially more likely to take up insurance when made eligible, conditional on the same expected treatment effect heterogeneity in increase in healthcare utilization. (This is broadly consistent with health policy discussions regarding mistrust of the healthcare system).
Next we consider imposing treatment parity constraints on an unconstrained optimal policy (defined on these estimates). In Figure˜8 we display the objective value, and $\mathbb{E}[T(\pi)\mid A=a]$, for gender and race, respectively, enumerated over values of the constraint. We use costs of $2$ for the number of doctors visits and $1$ for enrollment in Medicaid (so $\mathbb{E}[T(\pi)\mid A=a]$ is on the scale of probability of enrollment). These costs were chosen arbitrarily. Finding optimal policies that improve disparities in group-conditional access can be done with relatively little impact to the overall objective value. These group-conditional access disparities can be reduced from $4$ percentage points ($0.04$) for gender and about $6$ percentage points $(0.06)$ for race at a cost of $0.01$ or $0.02$ in objective value (twice the number of doctors’ visits). On the other hand, relative improvements/compromises in access value for the “advantaged group" show different tradeoffs. Plotting the tradeoff curve for race shows that, consistent with the large differences in treatment responsivity we see for whites, improving access for blacks. Looking at this disparity curve given $\lambda$ however, we can also see that small values of $\lambda$ can have relatively large improvements in access for blacks before these improvements saturate, and larger $\lambda$ values lead to smaller increases in access for blacks vs. larger decreases in access for whites.
E.3 Additional Discussion, PSA-DMF case study